# Learning solution operator of dynamical systems with diffusion maps kernel ridge regression

### v1

@@ 4.1 Kernel Ridge Regression
In this section, we give a short overview of kernel ridge regression (KRR). In our context, given a set of training data $S=(X,Y)$, where $X=[{\bf x}_{1},\ldots,{\bf x}_{N}]$ and $Y=[{\bf y}_{1},\ldots,{\bf y}_{N}]$, the KRR is to find a map $\mathbf{f}:\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ of the following form,
$f_{j}=k(\cdot,X)\boldsymbol{\alpha}_{j}=\sum_{i=1}^{N}k(\cdot,{\bf x}_{i})\alpha_{j,i},$ (7)
that solves,
$\min_{\boldsymbol{\alpha}_{j}}\sum_{i=1}^{N}\left((Y_{ji}-f_{j}({\bf x}_{i})\right)^{2}+{\lambda_{\mathrm{reg}}}\|\boldsymbol{\alpha}_{j}\|^{2},$ (8)
for each component $j=1\ldots,N$, where $\lambda_{\mathrm{reg}}$} is a user-specified ridge parameter for regularization. In the notation above, we denote $Y_{ji}$ as the $(j,i)$th component of $Y$. This problem admits a unique solution with $\boldsymbol{\alpha}_{j}=(\mathbf{K}_{N}+{\lambda_{\mathrm{reg}}}\mathbf{I}_{N})^{-1}Y_{j}^{\top}$, where $\mathbf{K}_{N}$ is an $N\times N$ Gram matrix with $(i,j)$-th component $k({\bf x}_{i},{\bf x}_{j})$, $\mathbf{I}_{N}$ is an identity matrix of size $N\times N$, and $Y_{j}$ denotes the $j$th row of data matrix $Y$.
@@
While any smooth kernel can be used, we have used the following Gaussian radial basis function (RBF) kernel as a baseline method in our numerical simulation in previous section,
$\tilde{k}_{\epsilon}({\bf x},{\bf y})=\exp\left(-\frac{\|{\bf x}-{\bf y}\|^{2}}{4\epsilon}\right),$ (9)
where $\epsilon>0$ is the lengthscale hyperparameter and needs to be specified by the user. In the next section, we will introduce the Diffusion Maps Kernel.
4.2 Diffusion Maps Kernel
@@
A scalar-valued kernel that is found to be effective (as shown in our numerical results) is induced by the Diffusion Maps (DM) algorithm [coifman2006diffusion]. The algorithm constructs a transition matrix based on a Markov chain representation of heat diffusion on the data points. Closer points are more likely to transition to each other, and the transition matrix is constructed to reflect this. This algorithm has multiple possible normalizations to remove biases induced by the sampling distribution. Given a dataset $X=\{{\bf x}_{1},\ldots,{\bf x}_{N}\}$ (drawn i.i.d. with sampling measure $\mu$) lies on a $d$-dimensional, smooth, compact, Riemannian submanifold of $\mathcal{M}\subset\mathbb{R}^{n}$, our chosen normalization corresponds to a Graph Laplacian matrix that converges pointwisely and spectrally to the semigroup solution of the Laplace-Beltrami operator [coifman2005geometric, coifman2006diffusion, berry2016variable, dunson2021spectral, calder2022lipschitz, Peoples2025]. In this case, the eigenfunctions of the Laplace-Beltrami operator form a complete basis for the integrable functions on the manifold, and hence are arguably effective in representing the solution operators in this study.
@@
For simplicity of discussion, we only consider the fixed bandwidth diffusion maps algorithm with the appropriate normalizations to generate our desired Laplace-Beltrami operator; although a symmetric kernel induced by the variable bandwidth diffusion maps [berry2016variable] or the symmetrized kernel in [Peoples2025] could also be used in principle. To construct our kernel, we begin with the symmetrically weighted Gaussian RBF, $\tilde{k}_{\epsilon}({\bf x},{\bf y})$, as defined in (9) and then perform two normalization stages. The “right” normalization in DM is to remove the bias induced by non-uniform distribution of the data and is defined as,
@@
where $q_{\epsilon}({\bf x})=\int_{\mathcal{M}}\tilde{k}_{\epsilon}({\bf x},{\bf y})q({\bf y})\,d\text{Vol}({\bf y})$, with $q$ denoting the sampling density of the dataset ($d\mu=q\text{Vol}$). We note that $\hat{k}_{\epsilon}$ is symmetric and positive definite. The second normalization is with respect to $\hat{q}_{\epsilon}({\bf x})=\int_{\mathcal{M}}\hat{k}_{\epsilon}({\bf x},{\bf y})q({\bf y})d\text{Vol}({\bf y})$ to attain the Markov transition kernel,
@@
In SI A.6, we characterize the function space, or formally, the RKHS corresponding to this kernel corresponding to the DM kernel in (12). When the forward invariant set $\mathcal{M}$ is a smooth Riemannian manifold, we give a conjecture of the limiting RKHS as $N\to\infty$ and $\epsilon\to 0$ based on recent results in the literature.
4.3 Selection of Hyperparameters
The KRR formulation considered in this study has two hyperparameters: the regularization factor, $\lambda_{\mathrm{reg}}$}, in the KRR, and the lengthscale hyperparameter, $\epsilon$, in both the RBF and DM kernels. The choice of the two hyperparameters critically impacts the accuracy of the learned solution operator. Essentially, the second key finding of this study is that the validation procedure is instrumental in identifying the best hyperparameters to use, which we explain below.
Suppose we have a training dataset, following the form given in Section 4.1, and a validation dataset consisting of a number of trajectories. Then, a random search strategy is used for the validation procedure. It starts with a given range of hyperparameters $H=[\epsilon_{l},\epsilon_{h}]\times[{\lambda_{\mathrm{reg}}}_{,l},{\lambda_{\mathrm{reg}}}_{,h}]$, a criterion of choice $E$ (e.g., RMSE for manifold case and VPT for chaotic case), and the maximum number of trials. For each trial,
  1. 1.
    Randomly select $(\epsilon,{\lambda_{\mathrm{reg}}})\in H$.
  2. 2.
    Use the training dataset to learn the solution operator, using $(\epsilon,{\lambda_{\mathrm{reg}}})$ for KRR.
  3. 3.
    Using the learned KRR model, predict on the validation trajectories, and compute the average criterion $\bar{E}$.
The model associated with the best criterion is used as the final model for testing.
The range of hyperparameters for random search may vary significantly among different problems. We used a heuristic choice of the range. First, a reference lengthscale value $\epsilon^{*}$ is computed using an estimation algorithm, provided in SI A.7, based on the input data for training. Then, using $\epsilon^{*}$ the symmetric RBF kernel matrix is constructed and suppose its minimum eigenvalue is $\lambda_{\min}$; choose the reference regularization factor ${\lambda_{\mathrm{reg}}}^{*}=|\lambda_{\min}|$. Subsequently, the range is chosen to be $H=[\epsilon^{*}\Delta\epsilon,\epsilon^{*}/\Delta\epsilon]\times[{\lambda_{\mathrm{reg}}}^{*},{\lambda_{\mathrm{reg}}}^{*}/\Delta{\lambda_{\mathrm{reg}}}]$, where, unless otherwise stated, $\Delta\epsilon=10^{-2},\Delta{\lambda_{\mathrm{reg}}}=10^{-4}$.
Competing interests. The authors declare no competing interests.
Appendix A Supplementary Information
A.1 Dynamics on Torus
Coordinate transformation.
$\theta=\arctan\frac{z_{2}}{z_{1}},\quad\phi=\arccos z_{3}.$
Validation.
The validation results are shown in Fig. 6. For each number of samples $N$, ambient dimension $n$, and the model, an independent validation study is performed using random search with a maximum of 4096 trials. The range of random search is determined using the approach provided in SI A.7. Black squares represent the reference hyperparameters found using the heuristic approach. One thing to notice is that as $N$ increases, reference $\lambda_{\mathrm{reg}}$} increases which is a typical behavior in the KRR models. Red cross marks represent the best hyperparameter combination found from the validation. The scattered dots represent the trials of the random search; if the RMSE of a trial is too large, the corresponding dot is not shown in the plot. It is noted that as $N$ increases, the black square coincides well with the red cross mark, which demonstrates the accuracy of the reference hyperparameters. In addition, interestingly this agreement is better for higher ambient dimension as at $N=2048$, they are already aligned well. The validation plots in the other examples are obtained based on the same procedure.
A.2 Lorenz-63 Dynamics
Validation.
The validation process uses $2N_{v}$ steps of trajectory, from which 3 overlapping segments of length $N_{v}$ are randomly sampled. The VPT with $\gamma=0.3$ is calculated for each segment of trajectory and averaged to be used for validation. The validation results for one typical training-validation data subset are shown in Fig. 8, where the hyperparameters are determined by maximizing the VPT over the 3 validation trajectories. The top row corresponds to the DM results, and the bottom row corresponds to the RBF results. Using the reference $\epsilon^{*}$, the optimal hyperparameters are found within the search bounds.
Sensitivity study on validation data size.
Figure 7 reports the mean VPT of all 250000 VPT’s as $N_{v}$ increases. After $N_{v}\geq 1500$, both DM and RBF do not show significant improvement in mean VPT, which justifies the choice of $N_{v}=1500$ for the other $N$’s. Also noteworthy is that, at lower $N_{v}$, specifically $N_{v}\leq 700$, the models show a longer VPT than the length of validation trajectory. Lastly, comparing the two KRR models, DM consistently outperforms RBF for all $N_{v}$, and the increasing trend in VPT plateaus later than RBF for both $N$; this indicates better representation capability of DM for this chaotic problem.
Distribution of test VPT.
Figure 9 represents the distributions of VPT of the DM and RBF models with $N=\{512,1024,2048,4096\}$, compared against the state-of-the-art random feature model, DeepSkip [mandal2025learning]. For a fair comparison, the statistics employed in [mandal2025learning] are reported. From the 500 models considered in Table 1, we test the models with 500 mutually disjoint test trajectories so that each model is only tested against a single test trajectory. The violin plot shows the distributions of the 500 VPT’s obtained from the 500 KRR models.
Figure 8: The validation results for Lorenz-63 dynamics. The scattered points correspond to individual trials from the random search; trials yielding NaN in the kernel coefficients are omitted from the visualization for clarity. Black squares denote hyperparameter values identified using the heuristic approach, while red crosses indicate the optimal hyperparameter combinations selected based on validation performance. Because the validation landscape exhibits multiple local maxima, one optimal combination is chosen arbitrarily among these maxima.
A.3 Kuramoto-Sivashinsky Equation, Chaotic Dynamics
Validation.
The validation results are shown in Fig. 10, where the hyperparameters are determined by maximizing the VPT ($\gamma=0.5$) over 3 validation trajectories of length $N_{v}$, which are randomly sampled from a trajectory of $2N_{v}$ steps. The values are averaged over 3 VPT’s obtained from the 3 validation trajectories. We consider reduced search range, $\Delta\epsilon=10^{-1},\Delta{\lambda_{\mathrm{reg}}}=10^{-2}$ while maintaining the same density as other cases, which yields $1024$ pairs of $(\epsilon,{\lambda_{\mathrm{reg}}})$.
Figure 10: The validation results for K-S chaotic dynamics. Black squares represent the hyperparameters found using the heuristic approach. Red cross marks represent the best hyperparameter combination found from the validation. The scattered dots represent the trials of the random search; if the RMSE of a trial is too large, the corresponding dot is not shown in the plot.
A.4 Kuramoto-Sivashinsky Equation, Travelling Dynamics
Data preprocessing.
[…]
Validation for the KRR models.
Hyperparameters of the baseline models.
A.5 Pitch-Plunge Flat Plate
Direct numerical simulation (DNS).
Figure 12: The computational domain of pitch-plunge flat plate.
Data visualization.
(a) $t=0$
(b) $t=1$
(c) $t=2$
(d) $t=3$
Dimension reduction.
The dataset size is reduced using the PCA to reduce the computational cost. The training data of the $12$ 200-step trajectories are collected as ${\bf X}\in\mathbb{R}^{n\times 2400}$, $n=179101$. The mean is $\bar{{\bf x}}=\frac{1}{2400}{\bf X}\mathbf{1}$, where $\mathbf{1}$ is a $2400$-dimension vector of $1$’s, and the mean is subtracted from the training data ${\bf X}^{\prime}={\bf X}-\bar{{\bf x}}\mathbf{1}^{\top}$. Next, a truncated singular value decomposition is applied, ${\bf X}^{\prime}\approx{\bf U}_{r}\boldsymbol{\Sigma}_{r}\mathbf{V}_{r}^{\top}$, where ${\bf U}_{r}\in\mathbb{R}^{n\times r},\boldsymbol{\Sigma}_{r}\in\mathbb{R}^{r\times r},{\bf V}_{r}\in\mathbb{R}^{2400\times r}$ and $n\ll r$. Then, a full-dimensional state $\bf x$} is reduced as $\tilde{{\bf x}}=\frac{1}{\sigma_{1}}{\bf U}_{r}^{\top}({\bf x}-\bar{\bf x})\in\mathbb{R}^{r}$, where $\sigma_{1}$ is the largest singular value and used to normalize the state.
The models are subsequently trained and perform prediction in the $r$-dimensional reduced state space. We choose $r=81$ for the DM and RBF models, which retains 99.199.1% energy of the data. More dimensions, $r=162$, are retained for ResDMD, to allow for the validation of truncation order in the ResDMD algorithm.
Validation for the KRR models.
The validation is performed using a weighted normalized RMSE (WNRMSE) criterion. For $J$ trajectories of length $T$ in the reduced space, $\{\tilde{{\bf x}}_{1:T}^{(j)}\}_{j=1}^{J}$, suppose the rollout prediction is $\{\hat{{\bf x}}_{1:T}^{(j)}\}_{j=1}^{J}$. The WNRMSE for $j$ is defined as,
$\text{WNRMSE}=\frac{\sqrt{\sum_{j=1}^{J}\sum_{i=1}^{T}\hat{w}_{i}^{2}\|\hat{{\bf x}}_{i}^{(j)}-\tilde{{\bf x}}_{i}^{(j)}\|_{2}^{2}}}{\sqrt{\sum_{j=1}^{J}\sum_{i=1}^{T}\|\tilde{{\bf x}}_{i}^{(j)}\|_{2}^{2}}},\quad\hat{w}_{i}=\frac{w_{i}}{\sqrt{\sum_{i=1}^{T}w_{i}^{2}}}$ (14)
For DM, the predictions diverge for a substantial portion of the search space, leading to the large empty region visible in Fig. 15; RBF also exhibits high errors in the same region. There is also a region on the top-right side that seems appropriate for both models, but the validation study did not select hyperparameters from this region for DM or RBF. Specifically, for DM, hyperparameters from this region produce predictions that decay over long horizons. For RBF, hyperparameters in this region do provide more accurate prediction than the one presented in the main text, however, they give decaying response in the long horizon and is still worse than the corresponding DM model. The validation results indicate that the lack of data, i.e., longer prediction horizon, poses challenge to validation for RBF, but less so for DM.
ResDMD.
(a) Full complex plane
(b) Zoomed in to the upper right side.
More test results.
(a) $t=3.0s$
(b) $t=42.2s$
(c) $t=77.5s$
A.6 Reproducing Kernel Hilbert Space Corresponding to the DM Kernel

### v2

@@ 2.1 Kernel Ridge Regression
Given a set of training data $S=(X,Y)$, where $X=[{\bf x}_{1},\ldots,{\bf x}_{N}]$ and $Y=[{\bf y}_{1},\ldots,{\bf y}_{N}]$, the KRR is to find a map $\mathbf{f}:\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ of the following form,
$f_{j}=k_{\epsilon}(\cdot,X)\boldsymbol{\alpha}_{j}=\sum_{i=1}^{N}k_{\epsilon}(\cdot,{\bf x}_{i})\alpha_{j,i},$ (3)
where $k_{\epsilon}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is a scalar-valued kernel. In the above notation, $k_{\epsilon}(\cdot,X):\mathbb{R}^{n}\to\mathbb{R}^{N}$, where we evaluate the kernel’s second input on all data points in $X$.
The classical kernel ridge regression (KRR) solution to this modeling problem is to solve
$\min_{\boldsymbol{\alpha}_{j}}L_{j}(\boldsymbol{\alpha}_{j}):=\min_{\boldsymbol{\alpha}_{j}}\sum_{i=1}^{N}\left((Y_{ji}-f_{j}({\bf x}_{i})\right)^{2}+{\lambda_{\mathrm{reg}}}\|\boldsymbol{\alpha}_{j}\|^{2},$ (4)
for each component $j=1\ldots,N$, where $\lambda_{\mathrm{reg}}$} is a user-specified ridge parameter for regularization. In the notation above, we denote $Y_{ji}$ as the $(j,i)$th component of $Y$. This problem admits a unique solution with $\boldsymbol{\alpha}_{j}=(\mathbf{K}_{N}+{\lambda_{\mathrm{reg}}}\mathbf{I}_{N})^{-1}Y_{j}^{\top}$, where $\mathbf{K}_{N}$ is an $N\times N$ Gram matrix with $(i,j)$-th component $k_{\epsilon}({\bf x}_{i},{\bf x}_{j})$, $\mathbf{I}_{N}$ is an identity matrix of size $N\times N$, and $Y_{j}$ denotes the $j$th row of data matrix $Y$.
@@
2.2 Dynamic-Aware Validation Strategy
The KRR formulation considered in this study has two hyperparameters: the regularization factor, $\lambda_{\mathrm{reg}}$}, in the KRR, and the lengthscale hyperparameter, $\epsilon$, in both the RBF and DM kernels. The choice of the two hyperparameters critically impacts the accuracy of the learned solution operator. The standard validation strategy is to find the optimal $\epsilon,{\lambda_{\mathrm{reg}}}$ over a range of hyperparameters $H=[\epsilon_{l},\epsilon_{h}]\times[{\lambda_{\mathrm{reg}}}_{,l},{\lambda_{\mathrm{reg}}}_{,h}]$, in the sense that they minimize the same loss function $L_{j}$ in (4) over a validation data set. Such a validation metric, while it is useful other application, is not appropriate for our purpose since we ultimately want to iterate $\mathbf{f}$ to predict the dynamical systems.
For our application, we require a validation dataset consisting of a number (usually about 3-12 in our numerical experiments) of trajectories of length $N_{v}$. Instead of minimizing the same loss function $L_{j}$, we consider the following validation metrics.
For a validation trajectory of length $N_{V}$, $\{{\bf x}_{i}\}_{i=1}^{N_{V}}$, suppose the rollout prediction is $\{\hat{{\bf x}}_{i}\}_{i=1}^{T}$, where $\hat{{\bf x}}_{i+1}={\bf f}(\hat{{\bf x}}_{i})$ for $i=1,2,\cdots,T-1$, with $\hat{{\bf x}}_{1}={\bf x}_{1}$ and $\bf{f}$ denoting the learned solution operator for a particular choice of $\epsilon,{\lambda_{\mathrm{reg}}}$. The validation metrics are defined as follows:
  1. 1.
    Root Mean Squared Error (RMSE): For dynamical systems with smooth forward invariant sets, we consider RMSE evaluated over the prediction horizon as
    $\text{RMSE}=\sqrt{\frac{1}{nN_{V}}\sum_{i=1}^{N_{V}}\left\|\hat{{\bf x}}_{i}-{\bf x}_{i}\right\|_{2}^{2}}.$ (5)
  2. 2.
    Valid Prediction Time (VPT): For chaotic dynamical systems, we consider the VPT, defined as the maximum duration for which the normalized trajectory error remains below a certain error tolerance $\gamma$,
    $\mathrm{VPT}=\frac{1}{T_{\Lambda}}\max\left\{i|E_{i}\leq\gamma\text{ for all }i\leq N_{V}\right\},$ (6)
Let $E$ be either RMSE or VPT as a validation metric. We conduct the following random search strategy for the validation procedure. It starts with a given range of hyperparameters $H=[\epsilon_{l},\epsilon_{h}]\times[{\lambda_{\mathrm{reg}}}_{,l},{\lambda_{\mathrm{reg}}}_{,h}]$, and the maximum number of trials. For each trial,
  1. 1.
    Randomly select $(\epsilon,{\lambda_{\mathrm{reg}}})\in H$.
  2. 2.
    Use the training dataset of size $N$ to learn the solution operator, using $(\epsilon,{\lambda_{\mathrm{reg}}})$ for KRR.
  3. 3.
    Using the learned KRR model, predict on the validation trajectories, and compute the average of the validation metric $E$, averaged over the number of validation trajectories.
The model associated with the smallest validation metric is used as the final model for testing.
The range of hyperparameters for random search may vary significantly among different problems. While it is possible to use a generic adaptive grid search approach, to reduce the computational cost, we used a heuristic choice of the range. First, a reference lengthscale value $\epsilon^{*}$ is computed using an estimation algorithm, provided in the Appendix A, based on the input data for training. Then, using $\epsilon^{*}$ the symmetric RBF kernel matrix is constructed and suppose its minimum eigenvalue is $\lambda_{\min}$; choose the reference regularization factor ${\lambda_{\mathrm{reg}}}^{*}=|\lambda_{\min}|$. Subsequently, the range is chosen to be $H=[\epsilon^{*}\Delta\epsilon,\epsilon^{*}/\Delta\epsilon]\times[{\lambda_{\mathrm{reg}}}^{*},{\lambda_{\mathrm{reg}}}^{*}/\Delta{\lambda_{\mathrm{reg}}}]$, where, unless otherwise stated, $\Delta\epsilon=10^{-2},\Delta{\lambda_{\mathrm{reg}}}=10^{-4}$.
3 Diffusion Maps Kernel
Although the KRR model above can be used with any kernel, the quality of the model prediction depends crucially on the choice of kernel being used. It is well known that choosing a kernel is typically problem dependent since each kernel is associated to a class of hypothesis models that may or may not be appropriate for some applications.
In Subsection 3.1, we introduce a data-driven scalar-valued kernel induced by the Diffusion Maps (DM) algorithm [coifman2006diffusion]. Besides the strong empirical performances across various examples to be reported in Section 4, we also provide a theoretical discussion in subsection 3.2 that hinges on a conjecture that may explain why the DM kernel is an appropriate choice for our applications.
3.1 Formulation of the DM Kernel
@@
The Diffusion Maps algorithm constructs a transition matrix based on a Markov chain representation of heat diffusion on the data points. Closer points are more likely to transition to each other, and the transition matrix is constructed to reflect this. This algorithm has multiple possible normalizations to remove biases induced by the sampling distribution. Given a dataset $X=\{{\bf x}_{1},\ldots,{\bf x}_{N}\}$ (drawn i.i.d. with sampling measure $\mu$) lies on a $d$-dimensional, smooth, compact, Riemannian submanifold of $\mathcal{M}\subset\mathbb{R}^{n}$, our chosen normalization corresponds to a Graph Laplacian matrix that converges pointwisely and spectrally to the semigroup solution of the Laplace-Beltrami operator [coifman2005geometric, coifman2006diffusion, berry2016variable, dunson2021spectral, calder2022lipschitz, Peoples2025]. In this case, the eigenfunctions of the Laplace-Beltrami operator form a complete basis for the integrable functions on the manifold, and hence are arguably effective in representing the solution operators in this study.
@@
For simplicity of discussion, we only consider the fixed bandwidth diffusion maps algorithm with the appropriate normalizations to generate our desired Laplace-Beltrami operator; although a symmetric kernel induced by the variable bandwidth diffusion maps [berry2016variable] or the symmetrized kernel in [Peoples2025] could also be used in principle. To construct our kernel, we perform two normalization stages, on the symmetrically weighted Gaussian RBF kernel,
$\tilde{k}_{\epsilon}({\bf x},{\bf y})=\exp\left(-\frac{\|{\bf x}-{\bf y}\|^{2}}{4\epsilon}\right).$ (7)
First, the “right” normalization in DM is to remove the bias induced by non-uniform distribution of the data and is defined as,
@@
where $q_{\epsilon}({\bf x})=\int_{\mathcal{M}}\tilde{k}_{\epsilon}({\bf x},{\bf y})q({\bf y})\,d\text{Vol}({\bf y})$, with $q$ denoting the sampling density of the dataset ($d\mu=qd\text{Vol}$). We note that $\hat{k}_{\epsilon}$ is symmetric and positive definite. The second normalization is with respect to $\hat{q}_{\epsilon}({\bf x})=\int_{\mathcal{M}}\hat{k}_{\epsilon}({\bf x},{\bf y})q({\bf y})d\text{Vol}({\bf y})$ to attain the Markov transition kernel,
@@
3.2 Reproducing Kernel Hilbert Space Corresponding to the DM Kernel